OpenStudy (anonymous):

5 (a + 1) = 2 (a + 6 ) solving each proportions

OpenStudy (anonymous):

5a+5=2a+12 5a-2a=12-5 3a=7 a=7/3

OpenStudy (anonymous):

you know it is not spuose to be a fraction

OpenStudy (anonymous):

2.3 repeating

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

7/3 = 2.3

OpenStudy (anonymous):

if you want 7/3 in decimal form just divide 7 by 3. it gives you 2.333333

OpenStudy (anonymous):

oh okay